OpenStudy (anonymous):

How do you simplify 75^1/2 - 27^1/2?

OpenStudy (anonymous):

\[\sqrt{75}+\sqrt{27}=\sqrt{25*3}+\sqrt{9*3}=5\sqrt{3}+3\sqrt{3}=8\sqrt{3}\] i hope its correct

OpenStudy (anonymous):

\[its 2\sqrt{3} sorry 5-3\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

\[5\sqrt{3}-3\sqrt{3}=2\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt(75) - \sqrt(27) = \sqrt(25)*\sqrt(3)-\sqrt(9)*\sqrt(3)\]\[=(\sqrt(25)-\sqrt(9))*\sqrt(3)= (5-3)*\sqrt(3)=2\sqrt(3)\]